Determine if the outcome is unusual. Consider as unusual any results that differs from the mean by more than 2 standard deviations. That is, unusual values are either less than μ- 2σ or greater than μ +2σ. According to AccuData Media Research, 36% of televisions within the Chicago city limits are tuned to "Eyewitness News" a 5:00 pm on Sunday nights. At 5:00 pm on a given Sunday, 2500 such televisions are randomly selected and checked to determine what is being watched. Would it be unusual to find that 919 of the 2500 televisions are turned to Eyewitness News? Yes or No and why?
Answer #1

Answer

given that p = 0.36

sample size is n = 2500

so, mean =n*p = 2500*0.36 = 900

and standard deviation = sqrt(n*p*(1-p))
= sqrt(2500*0.36*(1-0.36))

= 24

and unusual data is below and above 2 standard deviation

so, lower limit. = mean -2SD = 900-2(24) = 852

and upper limit = 900+2(24) = 948

therefore, 919 is not unusual because it is within 2 standard deviation limit